# The Trans-Neptunian Object (119951) 2002 KX14 as revealed by multiple stellar occultations

### v1

@@ 2.1.5 Occultation E - July 01, 2023
An angular diameter of 0.04 mas for the occulted star was estimated using Eq. 1 (giant star, see Fig. 2), with $V$ = 16.15 mag and $K$ = 11.90 mag from Zacharias et al. (2004). This corresponds to 0.65 km at 2002 KX14’s geocentric distance, $\Delta$, of 37.85 au. The Fresnel scale in this case is again 1.3 km. Since the cycle time for this observation was 0.5 seconds, while 2002 KX14’s velocity at the time of occultation was 23.74 km/s, this translates to a distance of 11.87 km. Given that this distance is orders of magnitude greater than that derived from the star’s angular size and Fresnel diffraction effects, stellar angular size and diffraction have a negligible impact on the derivation of ingress and egress times. Even so, they were taken into consideration.
@@ 3.1 Stellar occultations
The data were compiled and managed via the Tubitak Occultation Portal website (Kilic et al. 2022). Initially, CCD/CMOS image sequences were calibrated for bias and flat-field following standard procedures with the AstroImageJ (AIJ) software (Collins et al. 2017) when they were available and provided. For observations recorded in video format, we used the Planetary Imaging PreProcessor (PIPP [^3] https://pipp.software.informer.com) software to convert the video files to FITS format before proceeding with photometric analysis.
@@
We then conducted time series multi-aperture differential photometry in AIJ, measuring the flux of the occulted star (blended with 2002 KX14) and dividing by that of selected comparison stars within the field of view (FoV). This approach minimized systematic photometric errors arising from atmospheric fluctuations. As the event duration was only a few seconds, any flux variations due to the rotational variability of the object did not influence the resulting light curve. Aperture and sky background inner and outer radii were chosen to minimize noise in the data beyond the primary flux drop caused by the occultation. Error bars were derived from Poisson noise calculations and scaled to match the data’s standard deviation (see for details Ortiz et al. 2020). Finally, the ingress and egress timings, along with their uncertainties, were computed using the Python Stellar Occultation Reduction and Analysis (SORA) package (Gomes-Júnior et al. 2022). The resulting light curves are shown in the Fig. 4, 5, 6, 7, 8 for Occ. A, B, C, D, and E respectively. Table 4 provides the extracted ingress and egress times along with their associated uncertainties.
@@
To project the occultation chords onto the sky plane, we computed the astrometric right ascension and declination of 2002 KX14 relative to each observing site (ICRF frame, JPL##\##18 ephemeris) at ingress and egress times shown in Table 4. This was done using the JPL Horizons online service for solar system data and ephemeris calculations. Next, we subtracted these coordinates (the chord endpoints) from those of the occulted stars, then converted this difference to kilometers in the sky plane at 2002 KX14’s geocentric distance. Results for each occultation are displayed in Fig. 9. All plots are scaled uniformly (1100 $\times$ 1100 km) for consistent comparison. We note that the centroid position of Occ. E is shifted by approximately 1000 km along the x-axis compared to the others. This suggests an issue with the star’s right ascension. Indeed, we confirm that this star has a Renormalized Unit Weight Error (RUWE) of 1.50 in the Gaia catalog. RUWE quantifies the quality of a star’s astrometric fit relative to the expected value for a single, unperturbed source, with values above 1.4 indicating potential issues in the astrometric solution. However, this does not affect our results, as the objective of this study is to determine the shape of the object, which is independent of the star’s absolute position in our methodology.
@@
Occ. A includes 6 chords (12 endpoints), making it ideal for determining the most accurate projected ellipse. To fit the chord endpoints with an ellipse, we applied the numerically stable, non-iterative least-squares algorithm described by Halír̃ & Flusser (1998). The best-fit values are presented in Table 2. To account for ingress and egress time uncertainties in the fitted ellipse, we ran a Monte Carlo simulation, generating 1000 random endpoint clones based on a Gaussian distribution centered on each chord endpoint and scaled by the respective error. We then fitted an ellipse to each clone set, determining the final ellipse parameters by averaging the results from the 1000 fits, with the standard deviation representing the error. We chose 1000 clones as this quantity showed convergence, with no further improvement in precision from adding more clones. With these values, we obtain an average diameter of $392\pm 27$ km, lower than the thermal diameter of $455\pm 27$ km estimated by Vilenius et al. (2012).
@@
Notice that we acknowledge the existence of three additional chords uploaded to the Occult software repository (https://occultations.org/observing/software/occult/). However, we were unable to access the raw data necessary to ensure a consistent analysis alongside the other six chords, including the determination of ingress and egress times with their associated uncertainties using the same procedure, as well as accounting for information about the instruments or observatories. Consequently, this study includes only the six chords for which we have complete data. Nevertheless, we validated our solution by incorporating these three additional chords, simply plotting the user-provided ingress and egress times (without error bars), and also including the Stracise chord, which presents technical issues. We confirm that our fitted ellipse remains consistent with these four additional chords (Figure 11).
Table 2: Parameters of the fitted ellipse for the chords obtained on Occ. A
@@
Next, we determined the best-fitting ellipse for the remaining occultations. Since the number of chords is significantly lower, and particularly in the case of Occ. D or E, where only one or two chords are available ––to obtain the general equation of an ellipse, at least five points are required for a unique solution; otherwise, the solution degenerates––, we adopted a different approach. In these cases, we identified the best-fitting ellipse using a parameter grid search. To achieve this, we defined a chi-squared statistic that includes two components, a point-to-ellipse contribution ($A_{i}$), calculating the deviation of each observed point $(x_{i},y_{i})$ from an ellipse model and taking into account the uncertainties in the coordinates, and a component that penalizes deviations of the Occ. A ellipse parameters ($B$), weighted by their respective uncertainties. The reason for defining this penalty is that, as noted by Benecchi & Sheppard (2013), the object V-band magnitude exhibits rotational variations below 0.05 mag, meaning that the projected shape of 2002 KX14 cannot differ significantly from the one already determined throughout the rotation cycle. Moreover, since this object is located at geocentric distances greater than 37 AU, its viewing geometry has remained nearly unchanged. For instance, over the time span covered between the first and last occultation, the phase angle ranges from 0.23 to 1.45 degrees—a difference of just $\sim$ 1 degree. The same reasoning applies to the rest of parameters, such as the aspect angle, regardless of the rotation pole’s orientation. Eq. 3 summarizes the chi-squared statistic as follows:
@@
In Eq. 4, $x^{\prime}_{i}$ and $y^{\prime}_{i}$ are the coordinates in the rotated ellipsoidal reference frame (with $x_{0}$ and $y_{0}$ being the coordinates of the center of the ellipse) and $\sigma_{i}=\sqrt{\sigma_{x_{i}}^{2}+\sigma_{y_{i}}^{2}}$ is the combined uncertainty for each point. In Eq. 5, $u$ and $v$ are the semi-major and semi-minor axes, respectively; $\phi$ is the position angle; $u_{A}$, $v_{A}$, and $\phi_{A}$ are the Occ. A values for $u$, $v$, $\phi$, respectively; and $\sigma_{u}$, $\sigma_{v}$ and $\sigma_{\phi}$ their corresponding uncertainties. The best grid search solutions for Occ. B, C, D and E are summarized in Table 5. The diagrams for all occultations with their best-fitting ellipses are found in Fig. 10.
Finally, we subtract the centroids of each ellipse ($x_{0},y_{0}$) from their respective chords, centering them at the origin of the coordinate system. We have a total of 30 endpoints, and we repeat the least-squares fitting using the Monte Carlo method, as described for the Occ. A. This new fit (Fig. 1) provides a more precise determination of the semi-major and semi-minor axes, $u=241.0\pm 7.2$ and $v=157.1\pm 5.2$, leading to a final value for the average area-equivalent diameter of $389.2\pm 8.7$ km.
Figure 1: Chords of all the occultations centered in a common reference position as explained in the text, along with the best fitting ellipse. This ellipse has a semi-major axis ($v$) of $241.0\pm 7.2$ km, a semi-minor axis ($u$) of $157.1\pm 5.2$ km, and an average area-equivalent diameter of $389.2\pm 8.7$ km
@@ 3.2 Geometric Albedo
4 Conclusions and future work
@@
Objects with a variability lower than 0.15 mag are possibly Maclaurin spheroids (i.e., oblate bodies with two equal axes larger than the third), where this low variability is likely due to the presence of large-scale albedo features on their surface (Thirouin et al. 2010). Thus, given the minimal rotational variability of 2002 KX14, we can reasonably assume that the observed variations are due to albedo features. Since the observed axis ratio ($u/v$) is 1.53, if we assume a spheroidal shape and adopt a rotation period of 7 hours—consistent with the typical values estimated for trans-Neptunian objects (Duffard et al. 2009)—we derive a density of approximately 900 kg m-3 using the Chandrasekhar formalism (Chandrasekhar 1987). That said, it is important to note that this density estimate relies on several assumptions that are not directly supported by our dataset, such as a rotation period of 7 hours, although significantly longer periods are equally plausible.
The refined measurements of 2002 KX14’s physical properties demonstrate the effectiveness of stellar occultations in characterizing distant objects in the outer solar system. Currently, the projected shapes of only 13 other TNOs have been determined, and among them, 2002 KX14 will be the first cold classical object (although, as explained in Section 1, it is possible that it belongs to the hot population). Either way, regardless of its classification, these results contribute to the growing body of knowledge on the trans-Neptunian region.
During recent years, 2002 KX14 has been passing through the dense star fields of the Milky Way, complicating the acquisition of high-quality light curves necessary for determining its rotation period. This difficulty is further compounded by its low rotational amplitude, which remains under 0.05 mag. As the object moves away from this crowded region, opportunities will arise to collect high-precision photometric data. Combining such data with the results presented here will enable to determine a rotational period and a more comprehensive understanding of 2002 KX14’s physical properties. Additionally, further occultation campaigns may refine the shape and albedo estimated in this work, particularly if more multi-chord events are observed. These efforts could significantly reduce uncertainties in the current models, allowing for the generation of more accurate shape models and potentially determining the object’s spin axis orientation. Future studies will also integrate these results with thermal modeling and spectroscopy to better constrain the object’s surface composition and thermal properties, offering a deeper insight into its evolutionary history.
@@ Appendix B Stellar occultation analyses
Figure 7: Normalized flux values obtained from the two positive observations made during Occ. D on May 19, 2023. The solid line represents a square well fit to the observational data. The observing sites and the significance of the drop are indicated in the inserted labels.
Figure 8: Normalized flux values obtained from the two positive observations made during Occ. E on July 01, 2023. The solid line represents a square well fit to the observational data. The observing site and the significance of the drop are indicated in the inserted label.
@@
26 May, 2020 Marseilles (France) 0523’42.3”E 4318’21.1”N 68.9 114 mm WATEC910HX 2.0 1.0 NTP S/N low O. Schreurs
@@
555*For Millen-Bruch and Nandrin, time corrections of -2.5 seconds and -0.66 seconds, respectively, were applied. These cameras exhibit certain time offsets that depend on the observation mode (a detailed analysis can be found in the following link).
@@
Table 5: Parameters obtained through $\chi^{2}$ minimization for Occ. B, C, D and E.
@@
Figure 11: The six chords and the fitted ellipse for Occ. A on May 26, 2020, are shown alongside the three additional chords available in the Occult software. Since their precise locations are unknown, the legend displays the names of the users who uploaded the data. We also include the Stracise chord, which presents technical issues. The newly incorporated chords are represented as dashed-dotted lines in this figure. We confirm that they remain consistent with the values obtained in this study.

### v2

@@ 2.5 Occultation E: July 01, 2023
An angular diameter of 0.04 mas for the occulted star was estimated using Eq. 1 (giant star, see Fig. 2), with $V$ = 16.15 mag and $K$ = 11.90 mag from Zacharias et al. (2004). This corresponds to 0.65 km at 2002 KX14’s geocentric distance, $\Delta$, of 37.85 au. The Fresnel scale in this case is again 1.3 km. Since the cycle time for this observation was 0.5 seconds, while 2002 KX14’s velocity at the time of occultation was 23.74 km/s, this translates to a distance of 11.87 km. Given that this distance is several orders of magnitude greater than that derived from the star’s angular size and Fresnel diffraction effects, the stellar angular size and diffraction have a negligible impact on the derivation of ingress and egress times. Even so, they still were taken into consideration.
@@ 3.1 Stellar occultations
The data were compiled and managed via the Tubitak Occultation Portal website (Kilic et al. 2022). Initially, CCD/CMOS image sequences were calibrated for bias and flat-field following standard procedures with the AstroImageJ (AIJ) software (Collins et al. 2017) when they were available. For observations recorded in video format, we used the Planetary Imaging PreProcessor (PIPP [^4] https://pipp.software.informer.com) software to convert the video files to FITS format before proceeding with photometric analysis.
@@
We then conducted time series multi-aperture differential photometry in AIJ, measuring the flux of the occulted star (blended with 2002 KX14) and dividing by that of selected comparison stars within the field of view (FoV). This approach minimized systematic photometric errors arising from atmospheric fluctuations. As the event duration was only a few seconds, any flux variations due to the rotational variability of the object did not influence the resulting light curve. Aperture and sky background inner and outer radii were chosen to minimize noise in the data beyond the primary flux drop caused by the occultation. Error bars were derived from Poisson noise calculations and scaled to match the data’s standard deviation (see for details Ortiz et al. 2020). Finally, the ingress and egress timings, along with their uncertainties, were computed using the Python Stellar Occultation Reduction and Analysis (SORA) package (Gomes-Júnior et al. 2022). The resulting light curves are shown in the Fig. 4, 5, 6, and 7 for Occ. A, B, C, D, and E, respectively. Table 4 provides the extracted ingress and egress times along with their associated uncertainties.
@@
To project the occultation chords onto the sky plane, we computed the astrometric right ascension and declination of 2002 KX14 relative to each observing site (ICRF frame, JPL##\##18 ephemeris) at ingress and egress times shown in Table 4. This was done using the JPL Horizons online service for solar system data and ephemeris calculations. Next, we subtracted these coordinates (the chord endpoints) from those of the occulted stars, then converted this difference to kilometers in the sky plane at 2002 KX14’s geocentric distance. The results for each occultation are displayed in Fig. 8. All plots are scaled uniformly (1100 $\times$ 1100 km) for consistent comparison. We note that the centroid position of Occ. E is shifted by approximately 1000 km along the x-axis compared to the others. This suggests an issue with the star’s right ascension. Indeed, we confirm that this star has a renormalized unit weight error (RUWE) of 1.50 in the Gaia catalog. RUWE quantifies the quality of a star’s astrometric fit relative to the expected value for a single, unperturbed source, with values above 1.4 indicating potential issues in the astrometric solution. However, this did not affect our results, as the objective of this study is to determine the shape of the object, which is independent of the star’s absolute position with respect to our methodology.
@@
There are six chords (12 endpoints) in Occ. A, making it ideal for determining the most accurate projected ellipse. To fit the chord endpoints with an ellipse, we applied the numerically stable, non-iterative least-squares algorithm described by Halír̃ & Flusser (1998). The best-fit values are presented in Table 2. To account for ingress and egress time uncertainties in the fitted ellipse, we ran a Monte Carlo simulation, generating 1000 random endpoint clones based on a Gaussian distribution centered on each chord endpoint and scaled by the respective error. We then fitted an ellipse to each clone set, determining the final ellipse parameters by averaging the results from the 1000 fits, with the standard deviation representing the error. We chose 1000 clones as this quantity showed convergence, with no further improvement in precision from adding more clones. With these values, we obtain an average diameter of $392\pm 27$ km, lower than the thermal diameter of $455\pm 27$ km estimated by Vilenius et al. (2012).
@@
We notice that we acknowledge the existence of three additional chords uploaded to the Occult software repository [^5] https://occultations.org/observing/software/occult/. However, we were unable to access the raw data necessary to ensure a consistent analysis alongside the other six chords, including the determination of ingress and egress times with their associated uncertainties using the same procedure, as well as accounting for information about the instruments or observatories. Consequently, this study includes only the six chords for which we have complete data. Nevertheless, we were able to validate our solution by incorporating these three additional chords, simply plotting the user-provided ingress and egress times (without error bars), and also including the Stracise chord, which presents technical issues. We confirm that our fitted ellipse remains consistent with these four additional chords (Figure 10).
Table 2: Parameters of the fitted ellipse for the chords obtained on Occ. A.
@@
Next, we determined the best-fitting ellipse for the remaining occultations. For these other cases the number of chords is significantly lower, and particularly in the case of Occ. D or E, where only one or two chords are available. Given that at least five points are required for a unique solution, we adopted a different approach. In these cases, we identified the best-fitting ellipse using a parameter grid search. To achieve this, we defined a chi-squared statistic that includes two components: a point-to-ellipse contribution ($A_{i}$), calculating the deviation of each observed point $(x_{i},y_{i})$ from an ellipse model and taking into account the uncertainties in the coordinates, and a component that penalizes deviations of the Occ. A ellipse parameters ($B$), weighted by their respective uncertainties. The reason for defining this penalty is that, as noted by Benecchi & Sheppard (2013), the object V-band magnitude exhibits rotational variations below 0.05 mag, meaning that the projected shape of 2002 KX14 cannot differ significantly from the one already determined throughout the rotation cycle. Moreover, since this object is located at geocentric distances greater than 37 AU, its viewing geometry has remained nearly unchanged. For instance, over the time span covered between the first and last occultation, the phase angle ranges from 0.23 to 1.45 degrees, which is a difference of just $\sim$ 1 degree. The same reasoning applies to the rest of parameters, such as the aspect angle, regardless of the rotation pole’s orientation. Equation 3 summarizes the chi-squared statistic, as follows:
@@
In Eq. 4, $x^{\prime}_{i}$ and $y^{\prime}_{i}$ are the coordinates in the rotated ellipsoidal reference frame (with $x_{0}$ and $y_{0}$ being the coordinates of the center of the ellipse) and $\sigma_{i}=\sqrt{\sigma_{x_{i}}^{2}+\sigma_{y_{i}}^{2}}$ is the combined uncertainty for each point. In Eq. 5, $u$ and $v$ are the semi-major and semi-minor axes, respectively; $\phi$ is the position angle; $u_{A}$, $v_{A}$, and $\phi_{A}$ are the Occ. A values for $u$, $v$, and $\phi$, respectively; and $\sigma_{u}$, $\sigma_{v}$, and $\sigma_{\phi}$ are their corresponding uncertainties. The best grid search solutions for Occ. B, C, D, and E are summarized in Table 5. The diagrams for all occultations with their best-fitting ellipses are found in Fig. 9.
Finally, we subtracted the centroids of each ellipse ($x_{0},y_{0}$) from their respective chords, centering them at the origin of the coordinate system. This gave us a total of 30 endpoints and we repeated the least-squares fitting using the Monte Carlo method, as described for the case of Occ. A. This new fit (Fig. 1) provides a more precise determination of the semi-major and semi-minor axes, $u=241.0\pm 7.2$ and $v=157.1\pm 5.2$, leading to a final value for the average area-equivalent diameter of $389.2\pm 8.7$ km.
Figure 1: Chords of all the occultations centered in a common reference position, as explained in the text, along with the best fitting ellipse. This ellipse has a semi-major axis ($v$) of $241.0\pm 7.2$ km, a semi-minor axis ($u$) of $157.1\pm 5.2$ km, and an average area-equivalent diameter of $389.2\pm 8.7$ km
@@ 3.2 Geometric albedo
4 Conclusions and future works
@@
Objects with a variability lower than 0.15 mag are possibly Maclaurin spheroids (i.e., oblate bodies with two equal axes larger than the third), where this low variability is likely due to the presence of large-scale albedo features on their surface (Thirouin et al. 2010). Thus, given the minimal rotational variability of 2002 KX14, we can reasonably assume that the observed variations are due to albedo features. Since the observed axis ratio ($u/v$) is 1.53, if we assume a spheroidal shape and adopt a rotation period of 7 hours, which is consistent with the typical values estimated for trans-Neptunian objects (Duffard et al. 2009), we derive a density of approximately 900 kg m-3 using the Chandrasekhar formalism (Chandrasekhar 1987). That said, it is important to note that this density estimate relies on several assumptions that are not directly supported by our dataset, such as a rotation period of 7 hours, although significantly longer periods are equally plausible.
The refined measurements of 2002 KX14’s physical properties demonstrate the effectiveness of stellar occultations in characterizing distant objects in the outer solar system. Currently, the projected shapes of only 13 other TNOs have been determined; among them, 2002 KX14 will be the first cold classical object (although, as explained in Section 1, it is possible that it could belong to the hot population). Either way, regardless of its classification, these results contribute to the growing body of knowledge on the trans-Neptunian region.
In recent years, 2002 KX14 has been passing through the dense star fields of the Milky Way, complicating the acquisition of high-quality light curves needed for determining its rotation period. This difficulty is further compounded by its low rotational amplitude, which remains below 0.05 mag. As the object moves away from this crowded region, opportunities will arise to collect high-precision photometric data. Combining such data with the results presented here will enable the determination of a rotational period and a more comprehensive understanding of 2002 KX14’s physical properties. Additionally, further occultation campaigns may refine the shape and albedo estimated in this work, particularly if more multi-chord events are observed. These efforts could significantly reduce uncertainties in the current models, allowing for the generation of more accurate shape models and potentially determining the object’s spin axis orientation. Future studies will also integrate these results with thermal modeling and spectroscopy to better constrain the object’s surface composition and thermal properties, offering a deeper insight into its evolutionary history.
@@ Appendix B Stellar occultation analyses
Figure 7: Normalized flux values from the two positive detections, shown in (a) and (b), during Occ. D on May 19, 2023, while (c) presents the normalized flux from the positive detection during Occ. E on July 1, 2023. In all panels, the solid line represents a square-well fit to the observational data. The observing sites and the significance of the flux drop are indicated in the corresponding labels.
@@
26 May, 2020 Marseilles (France) 0523’42.3”E 4318’21.1”N 68.9 114 mm WATEC910HX 2.0 1.0 NTP S/N low Anonymous
@@
777*For Millen-Bruch and Nandrin, time corrections of -2.5 seconds and -0.66 seconds, respectively, were applied. These cameras exhibit certain time offsets that depend on the observation mode (a detailed analysis can be found in the following link).
@@
Table 5: Parameters obtained through $\chi^{2}$ minimization for Occ. B, C, D, and E.
@@
Figure 10: Six chords and the fitted ellipse for Occ. A on May 26, 2020, shown alongside the three additional chords available in the Occult software. Since the names of the locations are unknown, the legend displays the names of the users who uploaded the data. We also include the Stracise chord, which presents technical issues. The newly incorporated chords are represented as dashed-dotted lines in this figure. We confirm that they remain consistent with the values obtained in this study.